# AfriHate: A Multilingual Collection of Hate Speech and Abusive Language Datasets for African Languages

### v1

@@ Abstract
Shamsuddeen Hassan Muhammad1,2thanks: Equal contribution, Idris Abdulmumin3∗, Abinew Ali Ayele4,21, David Ifeoluwa Adelani5, Ibrahim Said Ahmad2,6, Saminu Mohammad Aliyu2, Nelson Odhiambo Onyango7, Lilian D. A. Wanzare7, Samuel Rutunda8, Lukman Jibril Aliyu9, Esubalew Alemneh10, Oumaima Hourrane11, Hagos Tesfahun Gebremichael4, Elyas Abdi Ismail10, Meriem Beloucif12, Ebrahim Chekol Jibril13, Andiswa Bukula14, Rooweither Mabuya14, Salomey Osei15, Abigail Oppong16, Tadesse Destaw Belay17,23, Tadesse Kebede Guge18, Tesfa Tegegne Asfaw4, Chiamaka Ijeoma Chukwuneke19, Paul Röttger20, Seid Muhie Yimam21, Nedjma Ousidhoum22 1Imperial College London, 2Bayero University Kano, 3DSFSI, University of Pretoria, 4Bahir Dar University, 5Mila, McGill University & Canada CIFAR AI Chair, 6Northeastern University, 7Maseno University, 8Digital Umuganda, 9HausaNLP, 10Haramaya University, 11Al Akhawayn University, 12Uppsala University, 13Istanbul Technical University, 14SADiLaR, 15University of Deusto, 16Independent Researcher, 17Instituto Politécnico Nacional, 18Addis Ababa University, 19Lancaster University, 20Bocconi University, 21University of Hamburg, 22Cardiff University, 23Wollo University Contact: s.muhammad@imperial.ac.uk, seid.muhie.yimam@uni-hamburg.de
@@ 3.1 Data Collection
To crowdsource additional keywords, we first asked native speakers to provide us with a list of hateful, abusive, or controversial keywords. To further diversify our lists, we contacted social media influencers, who asked their followers to share abusive or hateful keywords in their local languages by filling in a form created for anonymous collection. This strategy resulted in a broader vocabulary as the followers come from various backgrounds. Moreover, we used off-the-shelf curated hate speech lexicons from the PeaceTech Lab [^2] https://www.peacetechlab.org/the-peacetech-toolbox and the Hatebase. [^3] https://hatebase.org The lists were post-processed by native speakers prior to the collection of the tweets.

### v2

@@ Abstract
Shamsuddeen Hassan Muhammad1,2thanks: Equal contribution, Idris Abdulmumin3∗, Abinew Ali Ayele4,5, David Ifeoluwa Adelani6, Ibrahim Said Ahmad2,7, Saminu Mohammad Aliyu2, Nelson Odhiambo Onyango8, Lilian D. A. Wanzare8, Samuel Rutunda9, Lukman Jibril Aliyu10, Esubalew Alemneh11, Oumaima Hourrane12, Hagos Tesfahun Gebremichael4, Elyas Abdi Ismail11, Meriem Beloucif13, Ebrahim Chekol Jibril14, Andiswa Bukula15, Rooweither Mabuya15, Salomey Osei16, Abigail Oppong17, Tadesse Destaw Belay18,19, Tadesse Kebede Guge20, Tesfa Tegegne Asfaw4, Chiamaka Ijeoma Chukwuneke21, Paul Röttger22, Seid Muhie Yimam5, Nedjma Ousidhoum23 1Imperial College London, 2Bayero University Kano, 3DSFSI, University of Pretoria, 4Bahir Dar University, 5University of Hamburg, 6Mila, McGill University & Canada CIFAR AI Chair, 7Northeastern University, 8Maseno University, 9Digital Umuganda, 10HausaNLP, 11Haramaya University, 12Al Akhawayn University, 13Uppsala University, 14Istanbul Technical University, 15SADiLaR, 16University of Deusto, 17Independent Researcher, 18Instituto Politécnico Nacional, 19Wollo University, 20Addis Ababa University, 21Lancaster University, 22Bocconi University, 23Cardiff University Contact: s.muhammad@imperial.ac.uk, seid.muhie.yimam@uni-hamburg.de
@@ 3.1 Data Collection
To crowdsource additional keywords, we first asked native speakers to provide us with a list of hateful, abusive, or controversial keywords. To further diversify our lists, we contacted social media influencers, who asked their followers to share abusive or hateful keywords in their local languages by filling in a form created for anonymous collection. This strategy resulted in a broader vocabulary as the followers come from various backgrounds. We also used off-the-shelf curated hate speech lexicons from PeaceTech Lab [^2] https://www.peacetechlab.org/the-peacetech-toolbox and Hatebase. [^3] https://hatebase.org The lists were post-processed by native speakers prior to the collection of the tweets.